# Frequentist Validity of Epistemic Uncertainty Estimators

### v1

@@ B.4 Data Attribution
Experimental Details: We use the MNIST dataset. For training, we randomly select 600 images from each class to form the training dataset and keep the test dataset the same as the original.
@@
We use a two layer convolutional neural network with 3 fully connected layers. We use Adam (Kingma and Ba,, 2015) as an optimizer with a learning rate of 0.001, betas of 0.9 and 0.999, weight decay of 0.0001, and a batch size of 32. We train for 10 epochs.
@@
To calculate the influence function, we freeze all but the final two fully connected layers and calculate the hessian and gradients only with respect to the parameters in these layers. Specifically, as in Koh and Liang, (2017), the change in model parameters (for the final two layers) for an $\epsilon$ increase in the weighting of the $i^{th}$ training data point is $)-H_{\hat{\theta}}^{-1}\nabla_{\theta}\log(\hat{p}_{Y_{i}}(X_{i};\hat{\theta}$)) where $\hat{\theta}$ is the MLE on the original (unweighted) dataset, $)H_{\hat{\theta}}=\frac{1}{n}\sum_{i=1}^{n}\nabla^{2}_{\theta}\log(\hat{p}_{Y_{i}}(X_{i};\hat{\theta}$)) is the hessian and $\nabla_{\theta}$ denotes derivative with respect to $\theta$. We use multinomial bootstrap weights. For practical calculation, we add the identity matrix scaled by $10^{-5}$ to the Hessian to ensure invertibility in cases with non-negative definite hessian—see Koh and Liang, (2017) for discussion.
@@ B.4.1 Datamodels Experiment
We now also investigate the use of datamodels (Ilyas et al.,, 2022) as a means of approximating our measure.
Experimental Details: We use the CIFAR-10 dataset (Krizhevsky,, 2009) and ResNet-9 architecture as in Ilyas et al., (2022). Each test point is associated with a datamodel which consists of a weight vector of length 50001—the number of training data points and a bias term. The datamodel is such that it predicts the model output for that test data point given the weights assigned to the training data as the dot product between the training data weights and the datamodel weights. The model output considered in Ilyas et al., (2022) is unfortunately not the entire probability vector, but simply the difference between the correct logit and the highest incorrect logit. For our experiment, we use the datamodels for CIFAR-10 provided as part of the paper—see Ilyas et al., (2022) for details. We then draw 100 sets of Dirichlet weights and obtain estimates of model outputs using the datamodel. We then pass the model output through the sigmoid function to obtain a probability and use this for calculation of the EU. Note that this is equivalent to treating the classification task as binary between the correct class vs the highest incorrect class. While this is not equivalent to the correct calculation of EU over the 10 class probability vector, it is the best we can do with the existing datamodel.
@@
Results: Fig. 4 shows that the datamodels approximation to our measure and the Dropout epistemic uncertainty scores have a high rank correlation and moderate Pearson correlation. The datamodel estimates are also consistently lower than the Dropout estimates. This may be because of the reduction to a binary classification task in the Datamodels approximation versus the full 10 class classification in the Dropout calculation. Nonetheless, this is a useful proof of concept experiment which opens the door to using existing data attribution methods for approximate calculation of our measure.

### v2

@@ B.2 Data Attribution
Experimental Details: We use the MNIST dataset (LeCun et al., 1998). For training, we randomly select 600 images from each class to form the training dataset and keep the test dataset the same as the original.
@@
We use a two layer convolutional neural network with 3 fully connected layers. We use Adam (Kingma and Ba, 2015) as an optimizer with a learning rate of 0.001, betas of 0.9 and 0.999, weight decay of 0.0001, and a batch size of 32. We train for 10 epochs.
@@
To calculate the influence function, we freeze all but the final two fully connected layers and calculate the hessian and gradients only with respect to the parameters in these layers. Specifically, as in (Koh and Liang, 2017), the change in model parameters (for the final two layers) for an $\epsilon$ increase in the weighting of the $i^{th}$ training data point is $)-H_{\hat{\theta}}^{-1}\nabla_{\theta}\log(\hat{p}_{Y_{i}}(X_{i};\hat{\theta}$)) where $\hat{\theta}$ is the MLE on the original (unweighted) dataset, $)H_{\hat{\theta}}=\frac{1}{n}\sum_{i=1}^{n}\nabla^{2}_{\theta}\log(\hat{p}_{Y_{i}}(X_{i};\hat{\theta}$)) is the hessian and $\nabla_{\theta}$ denotes derivative with respect to $\theta$. We use multinomial bootstrap weights. For practical calculation, we add the identity matrix scaled by $10^{-5}$ to the Hessian to ensure invertibility in cases with non-negative definite hessian—see (Koh and Liang, 2017) for discussion.
@@ B.2.1 Datamodels Experiment
We now also investigate the use of datamodels (Ilyas et al., 2022) as a means of approximating our measure.
Experimental Details: We use the CIFAR-10 dataset (Krizhevsky, 2009) and ResNet-9 architecture as in (Ilyas et al., 2022). Each test point is associated with a datamodel which consists of a weight vector of length 50001—the number of training data points and a bias term. The datamodel is such that it predicts the model output for that test data point given the weights assigned to the training data as the dot product between the training data weights and the datamodel weights. The model output considered in (Ilyas et al., 2022) is unfortunately not the entire probability vector, but simply the difference between the correct logit and the highest incorrect logit. For our experiment, we use the datamodels for CIFAR-10 provided as part of the paper—see (Ilyas et al., 2022) for details. We then draw 100 sets of Dirichlet weights and obtain estimates of model outputs using the datamodel. We then pass the model output through the sigmoid function to obtain a probability and use this for calculation of the EU. Note that this is equivalent to treating the classification task as binary between the correct class vs the highest incorrect class. While this is not equivalent to the correct calculation of EU over the 10 class probability vector, it is the best we can do with the existing datamodel.
@@
Results: Fig. 5 shows that the datamodels approximation to our measure and the Dropout epistemic uncertainty scores have a high rank correlation and moderate Pearson correlation. The datamodel estimates are also consistently lower than the Dropout estimates. This may be because of the reduction to a binary classification task in the Datamodels approximation versus the full 10 class classification in the Dropout calculation. Nonetheless, this is a useful proof of concept experiment which opens the door to using data attribution methods for a more nuanced understanding of sources of uncertainty .